OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. x2 + 4x + y2 - 6y = -4

OpenStudy (perl):

is this the equation? $$\Large x^2 + 4x + y^2 -6y = -4 $$

OpenStudy (anonymous):

yes

OpenStudy (perl):

you can complete the square on x and y

OpenStudy (anonymous):

can you explain this to me?

OpenStudy (perl):

It is true that $$ \Large x^2 + bx + \left( \frac b 2 \right) ^2 = \left(x+ \frac b 2 \right)^2 $$

OpenStudy (perl):

can you verify this?

OpenStudy (anonymous):

I have to complete the square and I'm not sure how to do that

OpenStudy (perl):

that equation tells you how to complete the square. take the middle coefficient b, add (b/2) ^2

OpenStudy (anonymous):

the coefficient of when the like terms are grouped

OpenStudy (perl):

$$ x^2 + 4x + \color{red}{\left( \frac 4 2 \right)^2} + y^2 -6y+\color{blue}{\left( \frac {-6}{2} \right)^2} = -4+ \color{red}{\left( \frac 4 2 \right)^2}+ \color{blue}{\left( \frac {-6}{2} \right)^2} $$

OpenStudy (perl):

now use that identity above

OpenStudy (perl):

$$ x^2 + 4x + \color{red}{\left( \frac 4 2 \right)^2} + y^2 -6y+\color{blue}{\left( \frac {-6}{2} \right)^2} = -4+ \color{red}{\left( \frac 4 2 \right)^2}+ \color{blue}{\left( \frac {-6}{2} \right)^2} \\~\\ \large \left( x + \color{red}{\left( \frac 4 2 \right)}\right)^2 + \left( y +\color{blue}{\left( \frac {-6}{2} \right)}\right)^2 = -4+ \color{red}{\left( \frac 4 2 \right)^2}+ \color{blue}{\left( \frac {-6}{2} \right)^2} $$

OpenStudy (anonymous):

thank you i think i understand now